# Orientation double covers of non-orientable Lefschetz fibrations

### v1

@@ Abstract
commutes, where $id_{[0,1]}:[0,1]\to[0,1]$ is the identity map and $proj_{1}:[0,1]\times N_{g}\to[0,1]$ is the first projection. Since $\Phi$ is the trivialization map, the lower triangle of the diagram (4.4) is commutative. So, it follows that the diagram (4.4) commutes, which implies that $\widetilde{\Phi}$ is the trivialization map of $a^{*}(f\circ\pi)|_{(f\circ\pi)^{-1}(\Sigma-\Delta_{f})}$. From the commutativity of the diagram (4.4), we have $\widetilde{\Phi}(0,p)\in\{0\}\times\pi^{-1}(\pi(\widetilde{\Xi}^{-1}(p)))$ for all $p\in\Sigma_{g-1}$. Define a map $r:(f\circ\pi)^{-1}(y)\to(f\circ\pi)^{-1}(y)$ to be $r(x,+o_{x}):=(x,-o_{x})$ and $r(x,-o_{x}):=(x,+o_{x})$ for any $(x,+o_{x}),(x,-o_{x})\in(f\circ\pi)^{-1}(y)$, which is the orientation reversing involution belonging to $\mathrm{Aut}(\pi|_{(f\circ\pi)^{-1}(y)})$. Then for each $p\in\Sigma_{g-1}$, the set $\pi^{-1}(\pi(\widetilde{\Xi}^{-1}(p)))$ is rewritten as $\{\widetilde{\Xi}^{-1}(p),(r\circ\widetilde{\Xi}^{-1})(p)\}$. Noticing that $\widetilde{\Phi}$ is orientation preserving, we observe that $\widetilde{\Phi}(0,p)$ equals to $(0,\widetilde{\Xi}^{-1}(p))$ for all $p\in\Sigma_{g-1}$. Therefore, a monodromy representation $\rho_{f\circ\pi}:\pi_{1}(\Sigma-\Delta_{f\circ\pi},y)\to\mathcal{M}(\Sigma_{g-1})$ of $f\circ\pi$ can be defined as $\rho_{f\circ\pi}([a]):=[\widetilde{\Xi}\circ\widetilde{\Phi}_{1}]$ for all $a]\in\pi_{1}(\Sigma-\Delta_{f\circ\pi},y)$. Moreover, it immediately follows from the commutative diagram (4.4) that the following diagram is commutative:
@@
The above comuutative diagram (4.5) shows that $\widetilde{\Xi}\circ\widetilde{\Phi}_{1}$ is a lift of $\Xi\circ\Phi_{1}$, that is, $\eta([\widetilde{\Xi}\circ\widetilde{\Phi}_{1}])=[\Xi\circ\Phi_{1}]$ holds. This yields the assertion. ∎
The rest of this subsection, suppose that $\Sigma=\mathbb{D}^{2}$ or $\mathbb{S}^{2}$. Let $f:X\to\Sigma$ be a Lefschetz fibration of genus-$g$, and fix $y\in\Sigma-\Delta_{f}$ and a diffeomorphism $\Xi:f^{-1}(y)\to S_{g}$, here we asuume that $\Xi$ is orirentation preserving if $f$ is orientable. Also, let $\zeta_{1},\cdots,\zeta_{n}\in\{1,-1\}$. Choosing a Hurwitz system $(\mu_{1},\cdots,\mu_{n})$ of $f$, we obtain the vanishing cycle $C_{i}\subset S_{g}$ of $f$ determined by $\mu_{i}$ and the meridian loop $\alpha_{i}\in\pi_{1}(\Sigma-\Delta_{f},y)$ with respect to $\mu_{i}$ ($i=1,\cdots,n$). We first consider the case of $\Sigma=\mathbb{D}^{2}$. It is obvious that the product $\alpha_{1}\cdots\alpha_{n}$ is equal to the homotopy class $\partial\mathbb{D}^{2}]$ in $\pi_{1}(\mathbb{D}^{2}-\Delta_{f},y)$. Therefore, mapping each of both sides of this equation by a monodromy representation $\rho_{f}:\pi_{1}(\mathbb{D}^{2}-\Delta_{f},y)\to\mathcal{M}(S_{g})$ of $f$, the monodromy of $\partial\mathbb{D}^{2}]$ is represented as $t_{C_{1}}^{\zeta_{1}},\cdots,t_{C_{n}}^{\zeta_{n}}$. This product of Dehn twists is called a monodromy factorization of $f$. Next, consider the case of $\Sigma=\mathbb{S}^{2}$. Since $\pi_{1}(\mathbb{S}^{2}-\Delta_{f},y)=\langle\alpha_{1},\cdots,\alpha_{n}\thickspace|\thickspace\alpha_{1}\cdots\alpha_{n}=1\rangle$, the relation $\alpha_{1}\cdots\alpha_{n}=1$ is provided. We then get the relation $t_{C_{1}}^{\zeta_{1}},\cdots,t_{C_{n}}^{\zeta_{n}}=1$ in $\mathcal{M}(S_{g})$ by $\rho_{f}$, which is also said to be a monodromy factorization of $f$.
@@
For a given monodromy factorization $t_{{c}_{1};\theta_{{c}_{1}}}\cdots t_{{c}_{n};\theta_{{c}_{n}}}$ of $f$, we can get a Hurwitz system $(\mu_{1},\cdots,\mu_{n})$ such that each $c_{i}$ is the vanishing cycle of $f$ determined by $\mu_{i}$ ($i=1,\cdots,n)$. In particular, we get the meridian loop $\alpha_{i}\in\pi_{1}(\Sigma-\Delta_{f},y)$ with respect to $\mu_{i}$ $(i=1,\cdots n)$. Hence, by applying Proposition 4.2,
@@
Let $f_{1}:X_{1}\to\Sigma$ and $f_{2}:X_{2}\to\Sigma$ be Lefschetz fibrations. They are said to be isomorphic if there exist a diffeomorphism $H:X_{1}\to X_{2}$ and an orientation preserving diffeomorphism $h:\Sigma\to\Sigma$ such that $f_{2}\circ H=h\circ f_{1}$, here $H$ is orientation preserving if $f_{1}$ and $f_{2}$ are orientable.
@@
We remark that it follows from elementary transformations that the isomorphism class of an orientable Lefschetz fibration does not change if the following substitution is applied to its monodromy factorizations:
@@
for $\zeta_{i}\in\{1,-1\}\thickspace(i=1,\cdots,n)$ and disjoint simple closed curves $c_{i}$ and $c_{i+1}$ in $\Sigma_{g-1}$.
@@
which shows that $f\circ\pi$ is isomorphic to $f_{\mathrm{i-1}}\circ\pi_{\mathrm{i-1}}$ from Theorem 4.5. By the same augment, we also have that $f\circ\pi$ is isomorphic to $f_{\mathrm{i-2}}\circ\pi_{\mathrm{i-2}}$, where $f_{\mathrm{i-2}}:X_{\mathrm{i-2}}\to\Sigma$ is a genus-$g$ non-orientable Lefschetz fibration which is given by ${t_{c_{1};\theta_{c_{1}}}}\cdots{t_{c_{i-1};\theta_{c_{i-1}}}}({t_{c_{i+1};\theta_{c_{i+1}}}}{t_{c_{i+1};\theta_{c_{i+1}}}}^{-1}){t_{c_{i};\theta_{c_{i}}}}{t_{c_{i+1};\theta_{c_{i+1}}}}\cdots{t_{c_{n};\theta_{c_{n}}}}$ and $\pi_{\mathrm{i-2}}:\widetilde{X_{\mathrm{i-2}}}\to X_{\mathrm{i-2}}$ denotes the standard orientation double covering map. Moreover, let $f_{\mathrm{ii}}:X_{\mathrm{ii}}\to\Sigma$ be a genus-$g$ non-orientable Lefschetz fibration which is constructed by ${t_{\omega(c_{1});\omega_{*}(\theta_{c_{1}})}}\cdots{t_{\omega(c_{n});\omega_{*}(\theta_{c_{n}})}}$ for a fixed auto-diffeomorphism $\omega:N_{g}\to N_{g}$. It follows from the same discussion in the proof of Corollary 4.3 and the relation (4.3) that a monodromy factorization of $f_{\mathrm{ii}}\circ\pi_{\mathrm{ii}}$ is produced by $t_{\widetilde{\omega}(\gamma_{1})}t_{\widetilde{\omega}(\overline{\gamma_{1}})}^{-1}\cdots t_{\widetilde{\omega}(\gamma_{n})}t_{\widetilde{\omega}(\overline{\gamma_{n}})}^{-1}$, here $\pi_{\mathrm{ii}}:\widetilde{X_{\mathrm{ii}}}\to X_{\mathrm{ii}}$ is the standard orientation double covering map and $\widetilde{\omega}:\Sigma_{g-1}\to\Sigma_{g-1}$ is the orientation preserving lift of $\omega$. Theorem 4.5 thus states that $f\circ\pi$ is isomorphic to $f_{\mathrm{ii}}\circ\pi_{\mathrm{ii}}$. More generally, we can find that if monodromy factorizations of two genus-$g$ non-orientable Lefschetz fibrations $f_{i}:X_{i}\to\Sigma$ change each other by applying the transformations (i-1)-(ii) in Theorem 1.2 a finite number of times, then the genus-(g1)(g-1) achiral Lefschetz fibtations $f_{i}\circ\pi_{i}:\widetilde{X_{i}}\to\Sigma$ determined by the standard orientation double covering maps $\pi_{i}:\widetilde{X_{i}}\to X_{i}$ are isomorphic $(i=1,2)$. Hence, there are orientation preserving diffeomorphisms $\widetilde{H}:\widetilde{X_{1}}\to\widetilde{X_{2}}$ and $h:\Sigma\to\Sigma$ such that $(f_{2}\circ\pi_{2})\circ\widetilde{H}=h\circ(f_{1}\circ\pi_{1})$, i.e., we get the following commutative diagram:

### v2

@@ Abstract
commutes, where $id_{[0,1]}:[0,1]\to[0,1]$ is the identity map and $proj_{1}:[0,1]\times N_{g}\to[0,1]$ is the first projection. Since $\Phi$ is a trivialization map, the lower triangle of the diagram (4.4) is commutative. So, it follows that the diagram (4.4) commutes, which implies that $\widetilde{\Phi}$ is a trivialization map of $a^{*}(f\circ\pi)|_{(f\circ\pi)^{-1}(\Sigma-\Delta_{f})}$. From the commutativity of the diagram (4.4), we have $\widetilde{\Phi}(0,p)\in\{0\}\times\pi^{-1}(\pi(\widetilde{\Xi}^{-1}(p)))$ for all $p\in\Sigma_{g-1}$. Define a map $r:(f\circ\pi)^{-1}(y)\to(f\circ\pi)^{-1}(y)$ to be $r(x,+o_{x}):=(x,-o_{x})$ and $r(x,-o_{x}):=(x,+o_{x})$ for any $(x,+o_{x}),(x,-o_{x})\in(f\circ\pi)^{-1}(y)$, which is the orientation reversing involution belonging to $\mathrm{Aut}(\pi|_{(f\circ\pi)^{-1}(y)})$. Then for each $p\in\Sigma_{g-1}$, the set $\pi^{-1}(\pi(\widetilde{\Xi}^{-1}(p)))$ is equal to $\{\widetilde{\Xi}^{-1}(p),(r\circ\widetilde{\Xi}^{-1})(p)\}$. Noticing that $\widetilde{\Phi}$ is orientation preserving, we observe that $\widetilde{\Phi}(0,p)$ equals to $(0,\widetilde{\Xi}^{-1}(p))$ for all $p\in\Sigma_{g-1}$. Therefore, a monodromy representation $\rho_{f\circ\pi}:\pi_{1}(\Sigma-\Delta_{f\circ\pi},y)\to\mathcal{M}(\Sigma_{g-1})$ of $f\circ\pi$ can be defined as $\rho_{f\circ\pi}([a]):=[\widetilde{\Xi}\circ\widetilde{\Phi}_{1}]$ for all $a]\in\pi_{1}(\Sigma-\Delta_{f\circ\pi},y)$, where the map $\widetilde{\Phi}_{1}:\Sigma_{g-1}\to(f\circ\pi)^{-1}(y)$ is defined by $\widetilde{\Phi}_{1}(p):=(\widetilde{proj_{2}}\circ\widetilde{\Phi})(1,p)$ for all $p\in\Sigma_{g-1}$ and $\widetilde{proj_{2}}:[0,1]\times(f\circ\pi)^{-1}(\Sigma-\Delta_{f\circ\pi})\to(f\circ\pi)^{-1}(\Sigma-\Delta_{f\circ\pi})$ is the second projection. Moreover, it immediately follows from the commutative diagram (4.4) that the following diagram is commutative:
@@
The above comuutative diagram (4.5) shows that $\widetilde{\Xi}\circ\widetilde{\Phi}_{1}$ is a lift of $\Xi\circ\Phi_{1}$, that is, $\eta([\Xi\circ\Phi_{1}])=[\widetilde{\Xi}\circ\widetilde{\Phi}_{1}]$ holds. This yields the assertion. ∎
The rest of this subsection, suppose that $\Sigma$ equals to either $\mathbb{D}^{2}$ or $\mathbb{S}^{2}$. Let $f:X\to\Sigma$ be a Lefschetz fibration of genus-$g$, and fix $y\in\Sigma-\Delta_{f}$ and a diffeomorphism $\Xi:f^{-1}(y)\to S_{g}$, here we asuume that $\Xi$ is orirentation preserving if $f$ is orientable. Also, let $\zeta_{1},\cdots,\zeta_{n}\in\{1,-1\}$. Choosing a Hurwitz system $(\mu_{1},\cdots,\mu_{n})$ of $f$, we obtain the vanishing cycle $C_{i}\subset S_{g}$ of $f$ determined by $\mu_{i}$ and the meridian loop $\alpha_{i}\in\pi_{1}(\Sigma-\Delta_{f},y)$ of $f$ with respect to $\mu_{i}$ ($i=1,\cdots,n$). We first consider the case of $\Sigma=\mathbb{D}^{2}$. It is obvious that the product $\alpha_{1}\cdots\alpha_{n}$ is equal to the homotopy class $\partial\mathbb{D}^{2}]$ in $\pi_{1}(\mathbb{D}^{2}-\Delta_{f},y)$. Therefore, mapping both sides of this equation by a monodromy representation $\rho_{f}:\pi_{1}(\mathbb{D}^{2}-\Delta_{f},y)\to\mathcal{M}(S_{g})$ of $f$, the monodromy of $\partial\mathbb{D}^{2}]$ is represented as $t_{C_{1}}^{\zeta_{1}},\cdots,t_{C_{n}}^{\zeta_{n}}$. This product of Dehn twists is called a monodromy factorization of $f$. Next, consider the case of $\Sigma=\mathbb{S}^{2}$. Since $\pi_{1}(\mathbb{S}^{2}-\Delta_{f},y)=\langle\alpha_{1},\cdots,\alpha_{n}\thickspace|\thickspace\alpha_{1}\cdots\alpha_{n}=1\rangle$, the relation $\alpha_{1}\cdots\alpha_{n}=1$ is provided. We then get the relation $t_{C_{1}}^{\zeta_{1}},\cdots,t_{C_{n}}^{\zeta_{n}}=1$ in $\mathcal{M}(S_{g})$ by $\rho_{f}$, which is also said to be a monodromy factorization of $f$.
@@
We suppose that a Hurwitz system $(\mu_{1},\cdots,\mu_{n})$ of $f$ and the meridian loops $\alpha_{1},\cdots,\alpha_{n}$ about it give the mondromy factorization $t_{{c}_{1};\theta_{{c}_{1}}}\cdots t_{{c}_{n};\theta_{{c}_{n}}}$. Since $\Delta_{f\circ\pi}=\Delta_{f}$, we can take $(\mu_{1},\cdots,\mu_{n})$ as a Hurwitz system of $f\circ\pi$. In particular, it follows that $\alpha_{1},\cdots,\alpha_{n}$ are meridian loops of $f\circ\pi$. By applying Proposition 4.2,
@@
Let $f_{1}:X_{1}\to\Sigma$ and $f_{2}:X_{2}\to\Sigma$ be Lefschetz fibrations. They are said to be isomorphic if there exist a diffeomorphism $H:X_{1}\to X_{2}$ and an orientation preserving diffeomorphism $h:\Sigma\to\Sigma$ such that $f_{2}\circ H=h\circ f_{1}$, here $H$ is assumed to be orientation preserving if $f_{1}$ and $f_{2}$ are orientable.
@@
We remark that it follows from elementary transformations that the isomorphism class of an orientable Lefschetz fibration does not change if the following substitution is applied to its monodromy factorization:
@@
for $\zeta_{i}\in\{1,-1\}\thickspace(i=1,\cdots,n)$ and disjoint simple closed curves $c_{i}$ and $c_{i+1}$ in $\Sigma_{g}$.
@@
which shows that $f\circ\pi$ is isomorphic to $f_{\mathrm{i-1}}\circ\pi_{\mathrm{i-1}}$ from Theorem 4.5. By the same augment, we also have that $f\circ\pi$ is isomorphic to $f_{\mathrm{i-2}}\circ\pi_{\mathrm{i-2}}$, where $f_{\mathrm{i-2}}:X_{\mathrm{i-2}}\to\Sigma$ is a genus-$g$ non-orientable Lefschetz fibration which is given by ${t_{c_{1};\theta_{c_{1}}}}\cdots{t_{c_{i-1};\theta_{c_{i-1}}}}({t_{c_{i+1};\theta_{c_{i+1}}}}{t_{c_{i+1};\theta_{c_{i+1}}}}^{-1}){t_{c_{i};\theta_{c_{i}}}}{t_{c_{i+1};\theta_{c_{i+1}}}}\cdots{t_{c_{n};\theta_{c_{n}}}}$ and $\pi_{\mathrm{i-2}}:\widetilde{X_{\mathrm{i-2}}}\to X_{\mathrm{i-2}}$ denotes the standard orientation double covering map. Moreover, let $f_{\mathrm{ii}}:X_{\mathrm{ii}}\to\Sigma$ be a genus-$g$ non-orientable Lefschetz fibration which is constructed by ${t_{\omega(c_{1});\omega_{*}(\theta_{c_{1}})}}\cdots{t_{\omega(c_{n});\omega_{*}(\theta_{c_{n}})}}$ for a fixed auto-diffeomorphism $\omega:N_{g}\to N_{g}$. It follows from the same discussion in the proof of Corollary 4.3 and the relation (4.3) that a monodromy factorization of $f_{\mathrm{ii}}\circ\pi_{\mathrm{ii}}$ is produced by $t_{\widetilde{\omega}(\gamma_{1})}t_{\widetilde{\omega}(\overline{\gamma_{1}})}^{-1}\cdots t_{\widetilde{\omega}(\gamma_{n})}t_{\widetilde{\omega}(\overline{\gamma_{n}})}^{-1}$, here $\pi_{\mathrm{ii}}:\widetilde{X_{\mathrm{ii}}}\to X_{\mathrm{ii}}$ is the standard orientation double covering map and $\widetilde{\omega}:\Sigma_{g-1}\to\Sigma_{g-1}$ is the orientation preserving lift of $\omega$. Theorem 4.5 thus states that $f\circ\pi$ is isomorphic to $f_{\mathrm{ii}}\circ\pi_{\mathrm{ii}}$. More generally, we can find that if positive monodromy factorizations of two genus-$g$ non-orientable Lefschetz fibrations $f_{j}:X_{j}\to\Sigma$ change each other by applying the transformations (i-1)-(ii) in Theorem 1.2 a finite number of times, then the genus-(g1)(g-1) achiral Lefschetz fibtations $f_{j}\circ\pi_{j}:\widetilde{X_{j}}\to\Sigma$ determined by the standard orientation double covering maps $\pi_{j}:\widetilde{X_{j}}\to X_{j}$ are isomorphic $(j=1,2)$. Hence, there are orientation preserving diffeomorphisms $\widetilde{H}:\widetilde{X_{1}}\to\widetilde{X_{2}}$ and $h:\Sigma\to\Sigma$ such that $(f_{2}\circ\pi_{2})\circ\widetilde{H}=h\circ(f_{1}\circ\pi_{1})$, i.e., we get the following commutative diagram: